OpenStudy (anonymous):

PLEASE HELP! √(3x)-4=6 I know the answer is 300/6. I just need help explaining the steps. I just do it, but have trouble explaining.

OpenStudy (anonymous):

36 = 3x-4 40 = 3x x = 40 /3

OpenStudy (anonymous):

so what i didnt get, was even though the 3x was square root, you still divide. I understand you have to simplify.

OpenStudy (anonymous):

because 3x is like 3*x, so you do the opposite on the other side of the equation.

OpenStudy (amistre64):

sqrt(3x) -4 = 6 ; +4 each side sqrt(3x) = 10 ; ^2 both sides 3x = 100 ; /3 both sides x = 100/3

OpenStudy (anonymous):

ack... lol

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

I get too confused on simple mathematics. Im sorry. lol

OpenStudy (anonymous):

i don't sure the function is sqrt (3x) or sqrt (3x-4), thats why parentheses are important

OpenStudy (anonymous):

It is sqrt(3x) -4

OpenStudy (amistre64):

i dont get 300/6 tho...which is the 'right' answer lol

OpenStudy (anonymous):

Yeah, I dont know where I got 300/6

OpenStudy (amistre64):

probably off a sheet of paper you wadded up and threw in the trash where it belonged? :)

OpenStudy (anonymous):

Probably, I might have did 100/3 just typo maybe ? lol.

OpenStudy (anonymous):

I really appreciate the help, tremendously.

OpenStudy (amistre64):

you know; ive answered over 300 questions and still havent gotten the 1st answering machine badge? ...

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

How do I give a badge.... i would give it to you if I could lol. This is my first time using this site.

OpenStudy (amistre64):

badges are given thru performing certain tasks; not handed out really; medals for good answers can be given by the masses tho ;)

OpenStudy (anonymous):

I see, well I wish you the best of luck :) You'll definitely get it in time. I really appreciated the help. Seriously.

OpenStudy (anonymous):

brb

OpenStudy (amistre64):

you might have to hit f5 to refresh your browser to see any 'good answer' buttons next to our names tho

OpenStudy (anonymous):

Why is it, when i try it myself... now i am getting 4/3? I am just trying to do it and get the right answer :/

OpenStudy (anonymous):

nevermind, i see. 4+6. Ah. lol

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

now how would I write this out to double check? would it be.... \[\sqrt{3(100/3)}-4=6 then multiply both sides by 3?\]

OpenStudy (anonymous):

please disregard the one-wordness there lol

OpenStudy (amistre64):

add 4 to both sides first

OpenStudy (amistre64):

then square both sides

OpenStudy (anonymous):

i understand the method. I get confused with fractions

OpenStudy (anonymous):

ack....

OpenStudy (anonymous):

so wait....

OpenStudy (amistre64):

3*(100/3) = 100 sqrt(100 = 10)

OpenStudy (amistre64):

sqrt(100)=10 10-4 = 6 is true

OpenStudy (anonymous):

OOOOO

OpenStudy (anonymous):

wow...

OpenStudy (anonymous):

Wow... i feel stupid at the moment :/